# OMoE: Diversifying Mixture of Low-Rank Adaptation by Orthogonal Finetuning

### v1

@@ 1 Introduction
Large language models (LLMs) offer a versatile and task-agnostic foundation that underpins an extensive array of applications, including text generation Zhuang et al. (2023); Song et al. (2023), question answering Huang et al. (2024), reasoning Wang et al. (2024), and personalized chat-bots Achiam et al. (2023); Chan (2023). Due to the computational and memory resources required by full-parameter fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) methods have emerged Hu et al. (2021); Hayou et al. (2024); Liu et al. (2024); Kalajdzievski (2023); Li and Liang (2021). As a popular PEFT method, Low-Rank Adaptation (LoRA), is characterized by freezing the backbone parameters while lightweight low-rank parameter matrices are fine-tuned. LoRA marks a significant advancement in the practical use of LLMs.
@@
Figure 1: (a), (b) and (c) visualize representations of two activated experts using soft routing, Top-2 routing and our OMoE method, while Top-2 routing utilizes a total of 8 experts. The representations are obtained from layer 8 of LLaMA-2 7B by projecting the vectors into 2D space via t-SNE for qualitative analysis. The visualization of more layers is provided in Appendix F. (d), (e) and (f) illustrate performance comparison between soft routing, Top-2 routing and OMoE.
@@
Despite LoRA’s effectiveness, the issue of compromised quality in heterogeneous tasks motivates recent studies to explore the combination of mixture-of-experts (MoE) and LoRA for its modular design and versatility. MoE enables flexible and potentially more generalizable handling of diverse data within a single model by routing them to specialized experts Shi et al. (2024); Li et al. (2024); Liu et al. (2023b); Zhang et al. (2024); Tian et al. (2024); Lin et al. (2024); Chen et al. (2024). For more details, we provide a section for related works in Appendix A. Basically, for each input token, activated experts need to act synergistically to generate the output, which presents a compelling research question: How can activated experts collaborate efficiently to maximize the utilization of the model’s capacity?
@@
In practice, MoE stacks more experts and parameters to tackle complex problems, as it enables a broader spectrum of specialized knowledge. Recent works  Qian et al. (2024); Shi et al. (2024) indicate that merely raising the number of experts does not lead to significant improvements, and models’ performance approximately follows a logistic growth pattern. A direct hypothesis is that most experts collapse into similar distributions, thus lacking diversity and contributing minimally to overall performance. Our visualization of representation distributions in MoE under various routing strategies verifies this hypothesis. As shown in Fig.1 (a)(b), soft routing and top-2 routing in vanilla MoE exhibit apparent redundant areas in the representation space, indicating that experts redundantly learn the same feature, which leads to the underutilization of MoE’s modular design.
@@
MoE, similar to deep ensemble learning, has widely benefited from diversity to enhance robustness and efficiency Kim et al. (2024). However, existing MoE-based methods do not effectively address the issue of diversity, which is essential for enhancing models’ efficiency. Although selective updates in top-k routing promote dissimilarities among experts, it is not sufficient for significantly improving the diversity of experts. Current solutions promote diversity by introducing an additional objective as diversity booster, which alter the main learning objective Lan et al. (2024). In contrast, OMoE applys orthogonal constraints directly without compromising optimality, achieving diversity effectively. A qualitative illustration of OMoE is presented in Fig.1(c).
@@
Consequently, we propose a novel method named Orthogonal Mixture-of-Experts (OMoE), which is a resource-efficient MoE variant, to promote the diversity of experts. Specifically, OMoE orthogonalizes the representations generated by a mixture of experts via Gram-Schmidt precess, thus favoring dissimilarity and preserving the hyperspherical energy. The Gram-Schmidt precess applys orthogonal constraints directly to the architecture, thus OMoE keeps the learning objective unchanged, without compromising optimality. Mathematically, the manipulated orthogonal representations belong to particular Riemannian manifold where the inner producted is defined, known as Stiefel manifold Li et al. (2020). In this paper, we focus on fine-tuning MoE models to enhance diversity and improve the performance of LLMs. Our extensive experiments and analyses across diverse benchmarks demonstrate that OMoE consistently outperforms state-of-the-art PEFT baselines while significantly reducing tunable parameters by approximately 75%. Summary of our contributions:
@@
  • We systemically analysis that most experts collapse into similar distributions in MoE from the perspective of representations. And we show that promoting diversity is essential for achieving efficient PEFT.
@@
  • we propose a novel and resource-efficient MoE variant, named OMoE, that trains experts in an orthogonal manner to effectively promote diverity.
@@
  • Extensive expriments demonstrate that OMoE can consistently achieve stable performance while significantly reducing tunable parameters.
@@ 2.1 LoRA Basics
LoRA Hu et al. (2021) achieves performance comparable to fine-tuning on multiple benchmarks by introducing low-rank matrices into the model’s architecture, facilitating effective adaptation to specific tasks while keeping the original model weights $W_{0}$ frozen. A LoRA block consists of two sequential low-rank matrices $A$ and $B$ to fit the residual weights for fine-tuning. Then the updated forward computation is calculated through:
@@
, where $y\prime\in R^{d}$ is the output, $y$ is the original model’s output and $x\in R^{k}$ denotes the input token. $B\in R^{d×r},A\in R^{r×k}$ with rank $r\ll\text{min}(d,k)$. Normally, matrix $B$ is initialized with zeroes and matrix $A$ is initialized with Kaiming Uniform He et al. (2015) to force $\Delta~{}y=0$.
@@ 2.2 LoRA’s MoE Variants
MoE variants involes replacing the original LoRA modules with $n$ parallel minor LoRA submodules, referred as experts, denoted by $\{E_{i}=B_{i}A_{i}\}_{i=1}^{n}$. During training and inference, a router $g(\cdot;\mathbf{G})$, parameterized by $\mathbf{G}$, assigns the input token $x$ to a subset of LoRA experts, $g(\cdot;\mathbf{G})\in R^{n}$. Typically, the router $g$ consists of a simple linear layer followed by a softmax to normalize coefficient and then the forward computation is modified as:
@@
Here, $g_{i}(x;\mathbf{G})$ indicates the router’s output for the $i$-th expert and $g(x;\mathbf{G})$ can be designed as top-k routing, top-p routing, soft routing or top-1 switch routing.
@@ 2.3 MoE’s Dilemma: An Exploratory Analysis
Soft routing. Soft routing in mixture of experts (MoE) networks allows for a weighted average of inputs to be assigned to multiple experts, enabling more flexible and adaptive processing. This means $g(x;\mathbf{G})_{i}$ is nonzero for all experts.
Top-k routing. Top-k routing is a well-performed routing strategy of LLM and PEFT, which activates the $k$ experts with the highest values in $g(x;\mathbf{G})$. Then, the renormalized router value for the $i$-th expert, $\hat{g}(x;\mathbf{G})_{i}$, is defined as follows:
$\hat{g}_{i}(x)=\begin{cases}\frac{g_{i}(x)}{\sum_{j\in\operatorname{top}(\mathbf{w},2)}g_{j}(x)},&i\in\operatorname{top}(g(x),k)\\ 0,&i\notin\operatorname{top}(g(x),k)\end{cases}$ (3)
where $\operatorname{top}(g(x),k)$ return the indices of the largest $k$ elements in $g(x)$.
From the perspective of model performance, as illustrated in Fig.1(c)(d)(e), top-2 routing exhibits superior performance compared to soft routing. In terms of representation, we visualized representations of two activated experts using soft routing and top-2 routing, with the latter utilizing a total of eight experts. A qualitative illustration is presented in Fig. 1(a)(b). The representations in top-2 routing perform greater diversity compared to soft routing, suggesting the benefits of diversity. However, redundant area sticks in top-2 routing. Our analysis supports the conclusions of Chi et al. (2022) that if the hidden states are routed to activated experts, they will be pushed closer together. In this paper, we would like to force the representations more diverse, thus they can be more expressive and discriminative. A comprehensive analysis of Fig. 1, lead us to identify three key observations:
@@
Observation 3: Top-2 routing enhances the diversity of experts by selectively updating, which inevitably introduces a large number of experts. The issue of representation redundancy among these experts persists in Top-2 routing.
Building upon our observations, OMoE is proposed to enhance expert diversity and reduce representation redundancy. In this method, experts are trained orthogonally, with their representations residing within the Stiefel manifold. As illustrated in Fig. 1 (c), OMoE effectively reduces representation redundancy while significantly reducing tunable parameters. OMoE are trained end-to-end based on MoE and LoRA, which will be detailed in section  3. This method facilitates flexibility and applicability, even in diverse or heterogeneous training corpus.
@@ 3 OMoE: Orthogonal Mixture-of-Experts
Figure 2: The architecture of OMoE. OMoE consists of $n$ experts formed by multiple LoRAs combined with frozen transformer blocks. During the fine-tuning stage, expert representations are constrained to lie within the Stiefel manifold, thereby achieving orthogonalization.
In this section, we first introduce the basics of Stiefel manifold, and then elaborate on the architectural design of OMoE, as illustrated in Fig 2.
@@ 3.1 Optimization within Stiefel manifold
The Stiefel manifold, as a smooth and differentiable Riemannian manifold, has been widely used in neural network training Qiu et al. (2023); Hendawy et al. (2023); Liu et al. (2023a). In our context, we study Riemannian optimization for MoE in PEFT, which ensures that expert representations satisfy orthonormality constraints. A set of $k$ expert representations in $\mathbb{R}^{d}$ for a given token in LoRA can be transformed into orthonormal representations $E=[e_{1},...,e_{k}]\in\mathbb{R}^{d\times k}$ using the Gram-Schmidt process or Cayley transformation. These orthonormal representations belong to a topological space known as the Stiefel manifold, making MoE-based PEFT equivalent to Riemannian optimization on the Stiefel manifold.
@@
The Stiefel manifold $S(d,k)$ (where $d\geq k$), is a Riemannian manifold consisting of all $d\times k$ orthonormal matrices $S\in\mathbb{R}^{d\times k}$ such that $S^{T}S=I$. The Stiefel manifold $S(d,k)$ can be regarded as an embedded submanifold of a Euclidean space. Consequently, the Riemannian metric $\rho$ is defined using the Euclidean metric as: $\rho_{E}(Z_{1},Z_{2})=tr(Z_{1}^{T}Z_{2})$, where $Z_{1},Z_{2}$ are vectors in the tangent $T_{S}\mathcal{M}$.
We assume that MoE consists of a mixture of experts with learnable parameters $\phi=[\phi_{1},...,\phi_{k}]$, and their representations are formulated as $E_{\phi}=[e_{1},...,e_{k}]$. To ensure that the expert representations lie within the Stiefel manifold and promote diversity, we impose an orthogonality constraint on the optimization target of Supervised Fine-Tuning (SFT) Ouyang et al. (2022). This leads to the following formulation:
@@
where $I_{k}\in\mathbb{R}^{k\times k}$ is the identity matrix and $J(\phi)$ is the objective function for SFT. During the training process of $J(\phi)$, we ensure that the orthogonality constraint is satisfied by applying the Gram-Schmidt process to force orthogonalization.
@@
Table 1: The overall comparison of different PEFT methods for single-task learning, using base models with different number of parameters. Reported results are accuracy scores and most of the results are extracted from the original papers, and * indicates that the results are reproduced by running the provided source code. Bold and Underline indicate the best and the second-best results.
@@ 3.2 OMoE forward
$E^{\prime}=GS(E)$ (6)
where $W_{0}$ represent the frozen pretrained weights of the feedfoward network or attention layer. $g(h^{l})$ denotes the router employed to assign coefficients to specific experts for different hidden states, and $E^{\prime}$ represents the orthogonalized expert representations obtained after applying the Gram-Schmidt process, denoted as $GS(\cdot)$. Specifically, in $GS(\cdot)$, given the representations of experts $E=[e_{1},...,e_{k}]$, we use the Gram-Schmidt process to obtain a set of k-orthonormal representations of experts $E^{\prime}=[e_{1}^{\prime},...,e_{k}^{\prime}]$:
@@
The Gram-Schmidt process maps a set of linearly independent vector $E$ to a set of orthonormal vectors $E^{\prime}$. The representation of the $i$-th expert $e_{i}$ is projected in the orthogonal direction to the subspace spanned by the representation of all $i-1$ experts, which can satisfy the hard constraint in Eq. 4. Fig. 2 presents an illustration of how OMoE works. In this figure, the representations of experts are constrained to lie within the Stiefel manifold, resulting in a distinct difference in the directions of the vectors.
@@ 4 Experiments
In this section, we conduct a series of experiments and offer a succinct interpretation to evaluate OMoE.
@@ 4.1 Experiment Setting
Dataset and Benchmarks. we conduct experiments on a collection of tasks: (a) diverse commonsense reasoning datasets, including common-sense question-answer tasks, ARC-e and ARC-c Clark et al. (2018), OpenBookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), SocialIQA Sap et al. (2019), BoolQ Clark et al. (2019). (b) a science completion task, Hellaswag Zellers et al. (2019). (c) a fill-in-the-blank task, Winogrande Sakaguchi et al. (2021). We utilize the PEFT framework provided by  Hu et al. (2023); Li et al. (2024) for training on these datasets. We choose LLaMA-2 7B, LLaMA-3 8B and LLaMA-2 13B as our backbone model. The detailed statistics, and evaluation metrics can be found in Appendix B.
Baselines. We compare our method with the popular and well-performed baselines. First, we compare OMoE with LoRA, its variants and MoE variants: 1) LoRA Hu et al. (2021); 2) DoRA Liu et al. (2024); 3) MixLoRA Li et al. (2024); 4) AdaLoRA Zhang et al. (2023a); 5) MoELoRA Liu et al. (2023b); 6) MiLoRA Zhang et al. (2024). Then, we compare with other recent PEFT baselines: 1) Parallel-Adapter He et al. (2021); 2) Learned-Adapter Zhang et al. (2023b); 3) P-tuning v2 Liu et al. (2021); 4) IAPT Zhu et al. (2024); 5) BitFit Zaken et al. (2021); 6)$(\text{IA})^{3}$ Liu et al. (2022); 7) SSP Hu et al. (2022). Most of the results are extracted from the original papers, and some baselines are reproduced by running the provided source code. In the main results subsection, we choose LoRA, DoRA and MixLORA as representative baselines and the comparison of other baselines are provided in Appendix D.
Settings. To evaluate the effectiveness of our method, we apply it on the basis of LoRA and DoRA, which are labeled as OMoE-LoRA and OMoE-DoRA respectively in the experiment. Both OMoE-LORA and OMoE-LORA are configured with rank r = 16, incorporating 2 experts and a soft routing mechanism. When fine-tuning pretrained LLMs, we only consider the SFT setting. After receiving a prompt or instruction, all the predictions are generated using the language modeling output. Due to limited length, other experimental settings for the baseline methods and the training procedure are in Appendix C.

### v2

@@ 1 Introduction
Large language models (LLMs) offer a versatile and task-agnostic foundation that underpins an extensive array of applications, including text generation [47, 38], question answering [17], reasoning [40], and personalized chat-bots [1, 3]. Due to the computational and memory resources required by full-parameter fine-tuning, parameter-efficient fine-tuning (PEFT) methods have emerged [13, 9, 29, 18, 24]. As a popular PEFT method, low-rank adaptation (LoRA) is characterized by freezing the backbone parameters while lightweight low-rank parameter matrices are fine-tuned. LoRA marks a significant advancement in the practical use of LLMs.
@@
Figure 1: (a), (b) and (c) visualize representations of two activated experts using soft routing, Top-2 routing and our OMoE method, while Top-2 routing utilizes a total of 8 experts. The representations are obtained from layer 8 of LLaMA-2 7B by projecting the vectors into 2D space via t-SNE for qualitative analysis. The visualization of more layers is provided in Appendix D. (d), (e) and (f) illustrate performance comparison between soft routing, Top-2 routing and OMoE.
@@
Despite LoRA’s effectiveness, the issue of compromised quality in heterogeneous tasks motivates recent studies to explore the combination of mixture-of-experts (MoE) and LoRA for its modular design and versatility. MoE enables flexible and potentially more generalizable handling of diverse data within a single model by routing them to specialized experts [37, 22, 28, 43, 39, 25, 4]. For more details, we provide a section for related works in Section 5. In principle, for each input token, activated experts need to act synergistically to generate the output, which presents a compelling research question: How can activated experts collaborate efficiently to maximize the utilization of the model’s capacity?
@@
In practice, MoE stacks more experts and parameters to tackle complex problems, as it enables a broader spectrum of specialized knowledge. Recent works [33, 37] indicate that merely raising the number of experts does not lead to significant improvements, and models’ performance approximately follows a logistic growth pattern. A direct hypothesis is that most experts collapse into similar distributions, thus lacking diversity and contributing minimally to overall performance. Our visualization of representation distributions in MoE under various routing strategies verifies this hypothesis. As shown in Fig.1(a)(b), soft routing and top-2 routing in vanilla MoE exhibit apparent redundant areas in the representation space, indicating that experts redundantly learn the same feature, which leads to the underutilization of MoE’s modular design.
@@
MoE, similar to deep ensemble learning, has widely benefited from diversity to enhance robustness and efficiency [19]. However, existing MoE-based methods do not effectively address the issue of diversity, which is essential for enhancing models’ efficiency. Although selective updates in top-k routing promote dissimilarities among experts, it is not sufficient for significantly improving the diversity of experts. Current solutions promote diversity by introducing an additional objective as diversity booster, which alter the main learning objective [21, 8]. In contrast, the novelty of OMoE lies in its hard constraint instead of introducing regularization loss. OMoE applies orthogonal constraints directly without compromising optimality, achieving diversity effectively. A qualitative illustration of OMoE is presented in Fig. 1(c).
@@
Consequently, we propose a novel method named Orthogonal Mixture-of-Experts (OMoE), which is a resource-efficient MoE variant, to promote the diversity of experts. Specifically, OMoE orthogonalizes the representations generated by a mixture of experts via Gram-Schmidt process, thus favoring dissimilarity and preserving the hyperspherical energy. The Gram-Schmidt process applies orthogonal constraints directly to the architecture, thus OMoE keeps the learning objective unchanged, without compromising optimality. Mathematically, the manipulated orthogonal representations belong to particular Riemannian manifold where the inner product is defined, known as Stiefel manifold [23]. In this paper, we focus on fine-tuning MoE models to enhance diversity and improve the performance of LLMs. Our extensive experiments and analyses across diverse benchmarks demonstrate that OMoE consistently outperforms state-of-the-art PEFT baselines while significantly reducing tunable parameters by approximately 75%. Summary of our contributions:
@@
  • We systematically analyze that most experts collapse into similar distributions in MoE from the perspective of representations. And we show that promoting diversity is essential for achieving efficient PEFT.
@@
  • We propose a novel and resource-efficient MoE variant, named OMoE, that trains experts in an orthogonal manner to effectively promote diversity.
@@
  • Extensive experiments demonstrate that OMoE can consistently achieve stable performance while significantly reducing tunable parameters.
@@ 2.1 LoRA Basics
LoRA [13] achieves performance comparable to fine-tuning on multiple benchmarks by introducing low-rank matrices into the model’s architecture, facilitating effective adaptation to specific tasks while keeping the original model weights $W_{0}$ frozen. A LoRA block consists of two sequential low-rank matrices $A$ and $B$ to fit the residual weights for fine-tuning. The updated forward computation is calculated through:
@@
, where $y^{\prime}\in\mathbb{R}^{d}$ is the output, $y$ is the original model’s output and $x\in\mathbb{R}^{k}$ denotes the input token. $B\in\mathbb{R}^{d×r},A\in\mathbb{R}^{r×k}$ with rank $r\ll\min(d,k)$. Normally, matrix $B$ is initialized with zeroes and matrix $A$ is initialized with Kaiming Uniform [11] to force $\Delta y=0$.
@@ 2.2 LoRA’s MoE Variants
MoE variants involve replacing the original LoRA modules with $n$ parallel minor LoRA submodules, referred to as experts, denoted by $\{E_{i}=B_{i}A_{i}\}_{i=1}^{n}$. During training and inference, a router $g(\cdot;\mathbf{G})$, parameterized by $\mathbf{G}$, assigns the input token $x$ to a subset of LoRA experts, $g(\cdot;\mathbf{G})\in\mathbb{R}^{n}$. Typically, the router $g$ consists of a simple linear layer followed by a softmax to normalize coefficients and then the forward computation is modified as:
@@
Here, $g_{i}(x;\mathbf{G})$ indicates the router’s output for the $i$-th expert, and $g(x;\mathbf{G})$ can be designed as top-k routing, top-p routing, soft routing, or top-1 switch routing.
@@ 2.3 MoE’s Dilemma: An Exploratory Analysis
Soft routing. Soft routing in MoE allows for a weighted average of inputs to be assigned to multiple experts, enabling more flexible and adaptive processing. This means $g(x;\mathbf{G})_{i}$ is nonzero for all experts.
Top-k routing. Top-k routing is a well-performing routing strategy of LLM and PEFT, which activates the $k$ experts with the highest values in $g(x;\mathbf{G})$. Then, the renormalized router value for the $i$-th expert, $\hat{g}(x;\mathbf{G})_{i}$, is defined as follows:
$\hat{g}_{i}(x)=\begin{cases}\frac{g_{i}(x)}{\sum_{j\in\operatorname{top}(\mathbf{w},2)}g_{j}(x)},&i\in\operatorname{top}(g(x),k)\\ 0,&i\notin\operatorname{top}(g(x),k),\end{cases}$ (3)
where $\operatorname{top}(g(x),k)$ returns the indices of the largest $k$ elements in $g(x)$.
From the perspective of model performance, as illustrated in Fig.1(c)(d)(e), top-2 routing exhibits superior performance compared to soft routing. In terms of representation, we visualized representations of two activated experts using soft routing and top-2 routing, with the latter utilizing a total of eight experts. A qualitative illustration is presented in Fig. 1(a)(b). The representations in top-2 routing show greater diversity compared to soft routing, suggesting the benefits of diversity. However, redundant areas stick in top-2 routing. Our analysis supports the conclusions of [5] that if the hidden states are routed to activated experts, they will be pushed closer together. In this paper, we would like to force the representations to be more diverse, so that they can be more expressive and discriminative. A comprehensive analysis of Fig. 1, leads us to identify three key observations:
@@
Observation 3: Top-2 routing enhances the diversity of experts by selectively updating them, which inevitably introduces a large number of experts. The issue of representation redundancy among these experts persists in top-2 routing.
Building upon our observations, OMoE is proposed to enhance expert diversity and reduce representation redundancy. In our method, experts are trained orthogonally, with their representations residing within the Stiefel manifold. As illustrated in Fig. 1 (c), OMoE effectively reduces representation redundancy while significantly reducing tunable parameters. OMoE is trained end-to-end based on MoE and LoRA, which will be detailed in Section 3. This method facilitates flexibility and applicability, even in diverse or heterogeneous training corpus.
@@ 3 OMoE: Orthogonal Mixture-of-Experts
In this section, we first introduce the basics of the Stiefel manifold, and then elaborate on the architectural design of OMoE, as illustrated in Fig. 2.
Figure 2: The architecture of OMoE. During the fine-tuning stage, expert representations are constrained to lie within the Stiefel manifold, thereby achieving orthogonalization.
@@ 3.1 Optimization within Stiefel Manifold
The Stiefel manifold, as a smooth and differentiable Riemannian manifold, has been widely used in neural network training [34, 12, 26]. In our context, we study Riemannian optimization for MoE in PEFT, which ensures that expert representations satisfy orthonormality constraints. A set of $k$ expert representations in $\mathbb{R}^{d}$ for a given token in LoRA can be transformed into orthonormal representations $E=[e_{1},\ldots,e_{k}]\in\mathbb{R}^{d\times k}$ using the Gram-Schmidt process or Cayley transformation. These orthonormal representations belong to a topological space known as the Stiefel manifold, making MoE-based PEFT equivalent to Riemannian optimization on the Stiefel manifold.
@@
The Stiefel manifold $S(d,k)$ (where $d\geq k$) is a Riemannian manifold consisting of all $d\times k$ orthonormal matrices $S\in\mathbb{R}^{d\times k}$ such that $S^{T}S=I$. The Stiefel manifold $S(d,k)$ can be regarded as an embedded submanifold of a Euclidean space. Consequently, the Riemannian metric $\rho$ is defined using the Euclidean metric as: $\rho_{E}(Z_{1},Z_{2})=tr(Z_{1}^{T}Z_{2})$, where $Z_{1},Z_{2}$ are vectors in the tangent $T_{S}\mathcal{M}$.
We assume that MoE consists of a mixture of experts with learnable parameters $\phi=[\phi_{1},...,\phi_{k}]$, and their representations are formulated as $E_{\phi}=[e_{1},\ldots,e_{k}]$. To ensure that the expert representations lie within the Stiefel manifold and promote diversity, we impose an orthogonality constraint on the optimization target of supervised fine-tuning (SFT) [32]. This leads to the following formulation:
@@
where $I_{k}\in\mathbb{R}^{k\times k}$ is the identity matrix and $J(\phi)$ is the objective function for SFT. During the training process of $J(\phi)$, we ensure that the orthogonality constraint is satisfied by applying the Gram-Schmidt process to enforce orthogonalization.
@@
Table 1: The overall comparison of different PEFT methods for single-task learning, using base models with different numbers of parameters. Reported results are accuracy scores and most of the results are extracted from the original papers, and * indicates that the results are reproduced by running the provided source code. Bold and underline indicate the best and the second-best results.
@@ 3.2 OMoE Forward
$E^{\prime}=GS(E),$ (6)
where $W_{0}$ represents the frozen pretrained weights of the feedforward network or attention layer. $g(h^{l})$ denotes the router employed to assign coefficients to specific experts for different hidden states, and $E^{\prime}$ represents the orthogonalized expert representations obtained after applying the Gram-Schmidt process, denoted as $GS(\cdot)$. Specifically, in $GS(\cdot)$, given the representations of experts $E=[e_{1},\ldots,e_{k}]$, we use the Gram-Schmidt process to obtain a set of k-orthonormal representations of experts $E^{\prime}=[e_{1}^{\prime},\ldots,e_{k}^{\prime}]$:
@@
The Gram-Schmidt process maps a set of linearly independent vectors $E$ to a set of orthonormal vectors $E^{\prime}$. The representation of the $i$-th expert $e_{i}$ is projected in the orthogonal direction to the subspace spanned by the representation of all $i-1$ experts, which can satisfy the hard constraint in Eq. 4. Fig. 2 presents an illustration of how OMoE works. In this figure, the representations of experts are constrained to lie within the Stiefel manifold, resulting in a distinct difference in the directions of the vectors.
@@ 4 Experiments
In this section, we conduct a series of experiments and offer a succinct interpretation of the results to evaluate OMoE.
@@ 4.1 Experiment Setting
Datasets and Benchmarks. We conduct experiments on a collection of tasks: (a) Diverse commonsense reasoning datasets, including common-sense question-answering tasks, ARC-e and ARC-c [7], OpenBookQA [31], PIQA [2], SocialIQA [36], BoolQ [6]. (b) A science completion task, Hellaswag [42]. (c) A fill-in-the-blank task, Winogrande [35]. We utilize the PEFT framework provided by  [15, 22] for training on these datasets. We choose LLaMA-2 7B, LLaMA-3 8B and LLaMA-2 13B as our backbone models. The detailed statistics and evaluation metrics can be found in Appendix A.
Baselines. We compare our method with the popular and well-performing baselines. First, we compare OMoE with LoRA, its variants and MoE variants: 1) LoRA [13]; 2) DoRA [29]; 3) MixLoRA [22]; 4) AdaLoRA [44]; 5) MoELoRA [28]; 6) MiLoRA [43]. Then, we compare it with other recent PEFT baselines: 1) Parallel-Adapter [10]; 2) Learned-Adapter [45]; 3) P-tuning v2 [30]; 4) IAPT [46]; 5) BitFit [41]; 6)$(\text{IA})^{3}$ [27]; 7) SSP [14]. Most of the results are extracted from the original papers, and some baselines are reproduced by running the provided source code. In the main results subsection, we choose LoRA, DoRA, and MixLoRA as representative baselines and the comparison of other baselines is provided in Appendix C.
Settings. To evaluate the effectiveness of our method, we apply it on the basis of LoRA and DoRA, which are labeled as OMoE-LoRA and OMoE-DoRA respectively in the experiment. Both OMoE-LoRA and OMoE-DoRA are configured with rank r = 16, incorporating 2 experts and a soft routing mechanism. When fine-tuning pretrained LLMs, we only consider the SFT setting. After receiving a prompt or instruction, all the predictions are generated using the language modeling output. Due to space limitations, other experimental settings for the baseline methods and the training procedure are in Appendix B.
@@
Method ARC-e ARC-c BoolQ OBQA Avg.
MixLoRA 84.8 75.0 75.0 85.2 80.0
[…]
Table 3: Results on multi-task learning using the MixLoRA and OMoE. The backbone model is LLaMA-3 8B.
Rank $r$ ARC-e ARC-c BoolQ PIQA Average
2 73.4 50.5 72.8 82.4 69.8
4 75.3 51.6 73.7 83.7 71.1
8 77.2 52.1 73.1 83.8 71.6
16 79.3 56.6 73.5 84.5 73.5
32 79.7 57.7 72.5 82.8 73.2
Table 4: Accuracy comparison of OMoE-LoRA with varying ranks for LLaMA-2 7B on various benchmarks.